# Efficient 3D Perception on Multi-Sweep Point Cloud with Gumbel Spatial Pruning

### v1

@@ Abstract
This paper studies point cloud perception within outdoor environments. Existing methods face limitations in recognizing objects located at a distance or occluded, due to the sparse nature of outdoor point clouds. In this work, we observe a significant mitigation of this problem by accumulating multiple temporally consecutive LiDAR sweeps, resulting in a remarkable improvement in perception accuracy. However, the computation cost also increases, hindering previous approaches from utilizing a large number of LiDAR sweeps. To tackle this challenge, we find that a considerable portion of points in the accumulated point cloud is redundant, and discarding these points has minimal impact on perception accuracy. We introduce a simple yet effective Gumbel Spatial Pruning (GSP) layer that dynamically prunes points based on a learned end-to-end sampling. The GSP layer is decoupled from other network components and thus can be seamlessly integrated into existing point cloud network architectures. Without incurring additional computational overhead, we increase the number of LiDAR sweeps from 10, a common practice, to as many as 40. Consequently, there is a significant enhancement in perception performance. For instance, in nuScenes 3D object detection and BEV map segmentation tasks, our pruning strategy improves several 3D perception baseline methods.
@@ I INTRODUCTION
Figure 1: Sparse or missing outdoor point cloud data poses significant challenges for perception tasks, particularly at a distance (a1) or when objects are occluded (b1). Accumulating multiple temporally consecutive LiDAR sweeps enhances the density of the point cloud (a2) (b2), thereby facilitating perception. However, this improvement comes at the expense of increased computational cost. This work explores cost-effective strategies for leveraging multiple LiDAR sweeps.
The past several years have witnessed exponential growth in 3D perception techniques [1, 2, 3], including considerable point cloud-based methods [4, 5], primarily attributed to the rapid advancement and significant cost reduction of LiDAR sensors. In outdoor applications, such as autonomous driving, the utilization of LiDAR point cloud is essential for precise scene geometry measurement and dynamic object localization [6]. Typically, a LiDAR sensor is positioned atop the ego-car and performs a comprehensive [7] or partial [8] panoramic scan of the surrounding environment. However, this operational mechanism presents two inherent limitations. Firstly, the density of captured points exhibits a gradient, with higher density near the ego-car and sparser distribution at greater distances, thereby posing challenges for distant perception (Fig. 1 a1). Secondly, occluded objects remain undetected by the LiDAR sensor due to their limited or absent beam reflections (Fig. 1 b1). Consequently, these limitations significantly hinder the accuracy and reliability of LiDAR-based perception. To overcome these challenges, previous endeavors have introduced vision cues and pursued the fusion of data from multiple sensor modalities [9, 10, 11].
Nevertheless, even when considering point cloud data as the sole input, there exist approaches to mitigate the aforementioned issues. Autonomous vehicles are typically equipped with Inertial Measurement Units (IMUs) that can measure the pose of the ego-car. Leveraging this information, we can align and accumulate multiple temporally adjacent LiDAR sweeps into the current frame (Figure 1 a2 & b2). As a result, the sparse regions of the point cloud become denser, thereby facilitating the subsequent perception step. Despite the lack of a comprehensive investigation, this technique has been widely adopted as a common practice in point cloud-based tasks, such as object detection [12, 9, 13, 14] and semantic segmentation [15, 16, 17, 18] due to its consistent advantages.
@@
In light of the above consideration, we propose a simple yet efficient Gumbel Spatial Pruning (GSP) Layer to fulfill the dynamic pruning of LiDAR points. The GSP layer is a plug-and-play component that can be readily inserted after each computation layer. It learns a binary classifier through Gumbel Softmax to decide which points should be pruned. Our key finding is that it is crucial to use hard Gumbel Softmax which outputs discrete binary values rather than continuous logits. This way, we can use the discrete prediction for indexing and avoid thresholding the probability map during inference, keeping the data distribution consistent with training, and thus minimizing the performance loss brought by pruning. GSP layer uses neither additional supervision like pre-computed pruning label, nor prior cues such as feature magnitude [29]. The pruning is learned solely from the supervision of the task loss and a sparse regularization, in an end-to-end manner. Extensive experiments show that, by introducing the GSP layer, the sweep number can be enlarged to 4×4$\times$4 × without incurring additional computational overhead, either in terms of FLOPS or latency, and perception accuracy remarkably improves. In the nuScenes dataset, we improve several 3D perception networks, including the vanilla TransL [13] and LargeKernel3D [14], on both 3D object detection tasks and BEV map segmentation tasks. Compared with existing spatial pruning methods [29], we achieve a better trade-off between task performance and network efficiency.
@@ II Related Work
The point cloud is often utilized in the area of 3D perception tasks, carrying a great amount of spatial information. Recent researchers have focused on retrieving effective information from limited point cloud samples. In this section, we will introduce the up-to-date work on point cloud perception and point cloud pruning.
Single-sweep LiDAR perception. At the early stage of point cloud implementation, its usage often requires matching with RGB images to leverage the missing spatial information of a 2D input. In this case, a single sweep is merely an assistive input. Similarly, single sweep perception is also implemented in LIDAR-only perception tasks. Yan et al. [16] propose JS3C-Net for single sweep LiDAR point cloud segmentation via learning contextual shape priors. Barrera et al. [30] introduce BirdNet+ to deal with object detection with a single LiDAR sweep. However, these researches all fail to densify the target point cloud and have been gradually surpassed by multi-sweep methods in 3D perception tasks. Especially in outdoor scenes, a single sweep can hardly provide enough points to fully perceive the environment. For relatively small-scale target objects, the sparse point cloud would cause distortion or even absence. This weakness brings out the utility of multi-sweep perception.
Multi-sweep LiDAR perception. Multi-sweep perception was proposed because of the overwhelming point cloud information, especially in autonomous driving areas, where the combination of vehicle and LiDAR makes point cloud collection much easier. For datasets such as nuScenes [7] and Waymo [8], researchers tend to include multi-sweep point clouds in their perception modules and carry out various tasks. Lu et al. [21] propose a LinK-based backbone to solve detection and segmentation tasks. Multi-sweep perception effectively promotes the performance of various algorithms, but the increasing workload due to massive input leads to larger time and computing resource consumption. Liu et al. [26] propose MV-DeepSDF to reconstruct 3D vehicles from multi-sweep point clouds. Wang et al. [31] construct PointMotionNet which learns from a sequence of 3D LiDAR point clouds and performs well in multi-sweep semantic segmentation tasks. However, almost all mentioned measures are faced with increasing time and computing power consumption, compared with single-sweep algorithms.
@@ III Background and Motivation
Due to the unbounded nature of outdoor scenes, The point cloud obtained by a roof-mounted LiDAR sensor in autonomous vehicles, exhibits a gradient distribution, with higher point density in proximity to the ego-car and a sparser distribution at greater distances, thereby presenting challenges for distant perception. Besides, the LiDAR’s laser beams capture only the first intersecting surface along their direction, rendering occluded objects unperceivable. In prior arts [11, 13, 34, 16, 20], a common practice to mitigate this issue is to accumulate a history LiDAR sweeps into the current point cloud. Given the ego-poses at the current timestamp $t$ and a history timestamp $t-\Delta t$, together with the extrinsic parameters of the LiDAR sensor, we obtain the transformation matrix $P_{t-\Delta t,t}$ and then transform the history point cloud by $(\mathbf{x^{\prime}}_{t-\Delta t},1)^{T}=P_{t-\Delta t,t}(\mathbf{x}_{t-\Delta t},1)^{T}$, where $\mathbf{x}_{t-\Delta t}$ indicates the original spatial coordinate of a point and $\mathbf{x^{\prime}}_{t-\Delta t}$ is its aligned coordinate in the current timestamp. Multiple sweeps are then concatenated, forming an accumulated point cloud $\{\mathbf{x^{\prime}}_{t-\Delta t}\in\mathbf{X^{\prime}}_{t-\Delta t}|\Delta t=0,1,2,...,T\}$. To specify which timestep each point comes from, an additional temporal dimension is appended after the spatial coordinates. Consequently, we have 4D coordinates represented by $(\mathbf{x^{\prime}}_{t-\Delta t},\Delta t)^{T}$.
A typical choice of the sweep number is a mild $T=10$ as used in most previous LiDAR-based object detectors [35, 13, 11], which means a duration of 0.4 seconds for 25FPS streaming data, and the accumulated point cloud is of size around $10^{5}$. Intuitively, further enlarging $T$ may still bring additional improvements in perception performance, but also results in a dramatic increase in computation cost, and thus is well explored in existing works. In this paper, we validate the existence of these benefits by imposing more temporal sweeps beyond the common setting. More importantly, we introduce a new method that significantly reduces computation and enables us to enjoy such benefits at almost no cost.
@@ IV-A Dynamic Voxelization
In comparison to a single sweep of LiDAR point cloud, the accumulated point cloud exhibits significantly higher redundancy, especially in the proximity of the ego-car. To mitigate this redundancy, voxelization is commonly employed following established practices [1, 37, 35, 13]. Initially, the 3D space is partitioned into equidistant voxels, and the input points are grouped according to their voxel assignments. Subsequently, voxel features are encoded based on the points contained within each voxel. Existing methods typically adopt a hard voxelization [1] approach, which we have identified as having two limitations. Firstly, it imposes a predefined maximum point threshold per voxel, resulting in non-negligible information loss when excess points are rejected. Secondly, the implementation itself is inefficient, making it impractical to handle the increasing time consumption and computational cost when we accumulate more sweeps.
@@ V-A Experimental Setups
  • TransL: The baseline architecture we used, adopted from the lidar branch of TransFusion [13], which is composed of a VoxelNet [1] encoder, a SECOND [37] decoder, and a transformer head.
@@ V-C Ablation Studies
Generalization Ability. Besides the vanilla Transfusion method, we also implement the pruning strategy on LargeKernel3D [14], one of the SOTA LiDAR-based detection algorithms based on sparse convolution on the nuScenes dataset, in order to prove the generalization ability of our method. In Table. IV, we show the comparison between the LK3D baseline and the LK3D-GSP on the nuScenes 3D detection task. The proposed GSP provides a consistent reduction in GFLOPs and latency, with a $28.4\%/23.5\%/30.3\%$ reduction in latency. Moreover, the performance accuracy drops are kept at a relatively low level. It proves the generalization ability of our strategy compared to other baseline methods. Note that the experiments on LargeKernel3D are implemented on an NVIDIA 3090ti GPU.
@@
TABLE VIII: Hard v.s. Dynamic voxelization. The computation power consumption of the widely used hard voxelization grows prohibitively expensive when we use a large number of sweeps, i.e., 40. However, dynamic voxelization reduces the latency of the sparse backbone by 57×57$\times$57 ×, showing its necessity for the use of multi-sweep LiDAR points.
@@ 40 TransL Hard 8710 8622
Figure 5: Visualization of GSP in LiDAR view. We visualize the raw point clouds and the GSP-pruned point clouds from two scenes.
@@ VI Conclusion
In this paper, we observe the spatial sparsity in point cloud perception tasks and mitigate it by accumulating multiple LiDAR sweeps, achieving promising performance in multiple tasks, including 3D object detection and BEV map segmentation. We then investigate the ascending time and computation power consumption and propose Gumbel Spatial Pruning layer to alleviate such issues. Extensive experiments prove our strategy to be both effective and efficient on various perception tasks, compared with the vanilla backbone. We hope that our work will stimulate more related research in the future.

### v3

@@ Abstract
This paper studies point cloud perception within outdoor environments. Existing methods face limitations in recognizing objects located at a distance or occluded, due to the sparse nature of outdoor point clouds. In this work, we observe a significant mitigation of this problem by accumulating multiple temporally consecutive point cloud sweeps, resulting in a remarkable improvement in perception accuracy. However, the computation cost also increases, hindering previous approaches from utilizing a large number of point cloud sweeps. To tackle this challenge, we find that a considerable portion of points in the accumulated point cloud is redundant, and discarding these points has minimal impact on perception accuracy. We introduce a simple yet effective Gumbel Spatial Pruning (GSP) layer that dynamically prunes points based on a learned end-to-end sampling. The GSP layer is decoupled from other network components and thus can be seamlessly integrated into existing point cloud network architectures. Without incurring additional computational overhead, we increase the number of point cloud sweeps from 10, a common practice, to as many as 40. Consequently, there is a significant enhancement in perception performance. For instance, in nuScenes 3D object detection and BEV map segmentation tasks, our pruning strategy improves several 3D perception baseline methods.
@@ I INTRODUCTION
Figure 1: Sparse or missing outdoor point cloud data poses significant challenges for perception tasks, particularly at a distance (a1) or when objects are occluded (b1). Accumulating multiple temporally consecutive point cloud sweeps enhances the density of the point cloud (a2) (b2), thereby facilitating perception. However, this improvement comes at the expense of increased computational cost. This work explores cost-effective strategies for leveraging multiple point cloud sweeps.
The past several years have witnessed exponential growth in 3D perception techniques [1, 2, 3], including considerable point cloud-based methods [4, 5], primarily attributed to the rapid advancement and significant cost reduction of point cloud sensors. In outdoor applications, such as autonomous driving, the utilization of point cloud is essential for precise scene geometry measurement and dynamic object localization [6]. Typically, a point cloud sensor is positioned atop the ego-car and performs a comprehensive [7] or partial [8] panoramic scan of the surrounding environment. However, this operational mechanism presents two inherent limitations. Firstly, the density of captured points exhibits a gradient, with higher density near the ego-car and sparser distribution at greater distances, thereby posing challenges for distant perception (Fig. 1 a1). Secondly, occluded objects remain undetected by the point cloud sensor due to their limited or absent beam reflections (Fig. 1 b1). Consequently, these limitations significantly hinder the accuracy and reliability of point-cloud-based perception. To overcome these challenges, previous endeavors have introduced vision cues and pursued the fusion of data from multiple sensor modalities [9, 10, 11].
Nevertheless, even when considering point cloud data as the sole input, there exist approaches to mitigate the aforementioned issues. Autonomous vehicles are typically equipped with Inertial Measurement Units (IMUs) that can measure the pose of the ego-car. Leveraging this information, we can align and accumulate multiple temporally adjacent point cloud sweeps into the current frame (Figure 1 a2 & b2). As a result, the sparse regions of the point cloud become denser, thereby facilitating the subsequent perception step. Despite the lack of a comprehensive investigation, this technique has been widely adopted as a common practice in point cloud-based tasks, such as object detection [12, 9, 13, 14] and semantic segmentation [15, 16, 17, 18] due to its consistent advantages.
@@
In light of the above consideration, we propose a simple yet efficient Gumbel Spatial Pruning (GSP) Layer to fulfill the dynamic pruning of point clouds. The GSP layer is a plug-and-play component that can be readily inserted after each computation layer. It learns a binary classifier through Gumbel Softmax to decide which points should be pruned. Our key finding is that it is crucial to use hard Gumbel Softmax which outputs discrete binary values rather than continuous logits. This way, we can use the discrete prediction for indexing and avoid thresholding the probability map during inference, keeping the data distribution consistent with training, and thus minimizing the performance loss brought by pruning. GSP layer uses neither additional supervision like pre-computed pruning label, nor prior cues such as feature magnitude [29]. The pruning is learned solely from the supervision of the task loss and a sparse regularization, in an end-to-end manner. Extensive experiments show that, by introducing the GSP layer, the sweep number can be enlarged to 4×4$\times$4 × without incurring additional computational overhead, either in terms of FLOPS or latency, and perception accuracy remarkably improves. In the nuScenes dataset, we improve several 3D perception networks, including the vanilla TransL [13] and LargeKernel3D [14], on both 3D object detection tasks and BEV map segmentation tasks. Compared with existing spatial pruning methods [29], we achieve a better trade-off between task performance and network efficiency.
@@ II Related Work
The point cloud is often utilized in the area of 3D perception tasks, carrying a great amount of spatial information. Recent researchers have focused on retrieving effective information from limited point cloud samples [30, 31]. In this section, we will introduce the up-to-date work on point cloud perception and point cloud pruning.
Single-sweep point cloud perception. At the early stage of point cloud implementation, its usage often requires matching with RGB images to leverage the missing spatial information of a 2D input. In this case, a single sweep is merely an assistive input. Similarly, single sweep perception is also implemented in point-cloud-only perception tasks. Yan et al. [16] propose JS3C-Net for single sweep point cloud segmentation via learning contextual shape priors. Barrera et al. [32] introduce BirdNet+ to deal with object detection with a single point cloud sweep. However, these researches all fail to densify the target point cloud and have been gradually surpassed by multi-sweep methods in 3D perception tasks. Especially in outdoor scenes, a single sweep can hardly provide enough points to fully perceive the environment. For relatively small-scale target objects, the sparse point cloud would cause distortion or even absence. This weakness brings out the utility of multi-sweep perception.
Multi-sweep point cloud perception. Multi-sweep perception was proposed because of the overwhelming point cloud information, especially in autonomous driving areas, where the combination of vehicle and LiDAR makes point cloud collection much easier. For datasets such as nuScenes [7] and Waymo [8], researchers tend to include multi-sweep point clouds in their perception modules and carry out various tasks. Lu et al. [21] propose a LinK-based backbone to solve detection and segmentation tasks. Multi-sweep perception effectively promotes the performance of various algorithms, but the increasing workload due to massive input leads to larger time and computing resource consumption. Liu et al. [26] propose MV-DeepSDF to reconstruct 3D vehicles from multi-sweep point clouds. Wang et al. [33] construct PointMotionNet which learns from a sequence of 3D point clouds and performs well in multi-sweep semantic segmentation tasks. However, almost all mentioned measures are faced with increasing time and computing power consumption, compared with single-sweep algorithms.
@@ III Background and Motivation
Due to the unbounded nature of outdoor scenes, The point cloud obtained by a roof-mounted point cloud sensor in autonomous vehicles, exhibits a gradient distribution, with higher point density in proximity to the ego-car and a sparser distribution at greater distances, thereby presenting challenges for distant perception. Besides, the point cloud’s laser beams capture only the first intersecting surface along their direction, rendering occluded objects unperceivable. In prior arts [11, 13, 36, 16, 20], a common practice to mitigate this issue is to accumulate a history point cloud sweeps into the current point cloud. Given the ego-poses at the current timestamp $t$ and a history timestamp $t-\Delta t$, together with the extrinsic parameters of the point cloud sensor, we obtain the transformation matrix $P_{t-\Delta t,t}$ and then transform the history point cloud by $(\mathbf{x^{\prime}}_{t-\Delta t},1)^{T}=P_{t-\Delta t,t}(\mathbf{x}_{t-\Delta t},1)^{T}$, where $\mathbf{x}_{t-\Delta t}$ indicates the original spatial coordinate of a point and $\mathbf{x^{\prime}}_{t-\Delta t}$ is its aligned coordinate in the current timestamp. Multiple sweeps are then concatenated, forming an accumulated point cloud $\{\mathbf{x^{\prime}}_{t-\Delta t}\in\mathbf{X^{\prime}}_{t-\Delta t}|\Delta t=0,1,2,...,T\}$. To specify which timestep each point comes from, an additional temporal dimension is appended after the spatial coordinates. Consequently, we have 4D coordinates represented by $(\mathbf{x^{\prime}}_{t-\Delta t},\Delta t)^{T}$.
A typical choice of the sweep number is a mild $T=10$ as used in most previous point-cloud-based object detectors [37, 13, 11], which means a duration of 0.4 seconds for 25FPS streaming data, and the accumulated point cloud is of size around $10^{5}$. Intuitively, further enlarging $T$ may still bring additional improvements in perception performance, but also results in a dramatic increase in computation cost, and thus is well explored in existing works. In this paper, we validate the existence of these benefits by imposing more temporal sweeps beyond the common setting. More importantly, we introduce a new method that significantly reduces computation and enables us to enjoy such benefits at almost no cost.
@@ IV-A Dynamic Voxelization
In comparison to a single sweep of point cloud, the accumulated point cloud exhibits significantly higher redundancy, especially in the proximity of the ego-car. To mitigate this redundancy, voxelization is commonly employed following established practices [1, 39, 37, 13]. Initially, the 3D space is partitioned into equidistant voxels, and the input points are grouped according to their voxel assignments. Subsequently, voxel features are encoded based on the points contained within each voxel. Existing methods typically adopt a hard voxelization [1] approach, which we have identified as having two limitations. Firstly, it imposes a predefined maximum point threshold per voxel, resulting in non-negligible information loss when excess points are rejected. Secondly, the implementation itself is inefficient, making it impractical to handle the increasing time consumption and computational cost when we accumulate more sweeps.
@@ V-A Experimental Setups
  • TransL: The baseline architecture we used, adapted from the point cloud branch of TransFusion [13], which is composed of a VoxelNet [1] encoder, a SECOND [39] decoder, and a transformer head.
@@ V-C Ablation Studies
Generalization Ability. Besides the vanilla Transfusion method, we also implement the pruning strategy on LargeKernel3D [14], one of the SOTA point-cloud-based detection algorithms based on sparse convolution on the nuScenes dataset, in order to prove the generalization ability of our method. In Table. IV, we show the comparison between the LK3D baseline and the LK3D-GSP on the nuScenes 3D detection task. The proposed GSP provides a consistent reduction in GFLOPs and latency, with a $28.4\%/23.5\%/30.3\%$ reduction in latency. Moreover, the performance accuracy drops are kept at a relatively low level. It proves the generalization ability of our strategy compared to other baseline methods. Note that the experiments on LargeKernel3D are implemented on an NVIDIA 3090ti GPU.
@@
TABLE VIII: Hard v.s. Dynamic voxelization. The computation power consumption of the widely used hard voxelization grows prohibitively expensive when we use a large number of sweeps, i.e., 40. However, dynamic voxelization reduces the latency of the sparse backbone by 57×57$\times$57 ×, showing its necessity for the use of multi-sweep point cloud.
@@ 40 TransL Hard 8710 8622
Figure 5: Visualization of GSP in point cloud view. We visualize the raw point clouds and the GSP-pruned point clouds from two scenes.
@@ VI Conclusion
In this paper, we observe the spatial sparsity in point cloud perception tasks and mitigate it by accumulating multiple point cloud sweeps, achieving promising performance in multiple tasks, including 3D object detection and BEV map segmentation. We then investigate the ascending time and computation power consumption and propose Gumbel Spatial Pruning layer to alleviate such issues. Extensive experiments prove our strategy to be both effective and efficient on various perception tasks, compared with the vanilla backbone. We hope that our work will stimulate more related research in the future.